OpenStudy (anonymous):

The equation for the surface area of a cylinder is S=2piRh+2piR^2 slove for h

OpenStudy (anonymous):

\[S=2pirh+2pir^2\]

OpenStudy (anonymous):

is the radius given?

OpenStudy (anonymous):

no

OpenStudy (ghazi):

\[S-2*\pi*r^2=2*\pi*r*h\] then \[h= (S-2 \pi r^2)/2 \pi r\]

OpenStudy (anonymous):

Treat it like a linear equation to solve.

OpenStudy (ghazi):

is it clear?

OpenStudy (anonymous):

yea is tht all for it?